OpenStudy (unklerhaukus):

If a standard set of (52) playing cards is turned face up, how many diamonds are visible?

OpenStudy (kc_kennylau):

A-K: 2 diamonds each on top left hand corner and bottom right hand corner A-10: additional diamonds according to number So that's \(2\times13+1+2+3+4+5+6+7+8+9+10=2\times13+\dfrac{11\times10}2\)

OpenStudy (unklerhaukus):

=26+55 =81